Two liters of a 20% alcohol solution is mixed with x liters of

a 50% alcohol solution. if the new mixture contains 38%

alcohol, what is the value of x?


1
Expert's answer
2021-09-23T16:48:10-0400

Explanations & Calculations


  • The amount of alcohol provided to the final mixture by the 20% solution,

"\qquad\qquad
\begin{aligned}
\small q_1&=\small 2\times\frac{20}{100}\,l
\end{aligned}"

  • That from the 50% solution,

"\qquad\qquad
\begin{aligned}
\small q_2&=\small x\times\frac{50}{100}\,l
\end{aligned}"

  • These 2 amounts are included in the final mixture. The volume of the final mixture is,

"\qquad\qquad
\begin{aligned}
\small V&=\small (2+x)\,l
\end{aligned}"

  • Then from the final mixture,

"\qquad\qquad
\begin{aligned}
\small 38&=\small \Bigg[\frac{(2\times\frac{20}{100})+(x\times\frac{50}{100})}{2+x}\Bigg]\times100\\
\small x&=\small 3\,l
\end{aligned}"